A lorry of weight 15000N is parked on a metal bridge of length 10m at a distance of 2m from one end. If the weight of the bridge is 5000, what are the reactions at the two ends of the bridge?
1
Expert's answer
2020-03-09T10:55:21-0400

Let lorry is parked a distance of 2m from 1 end.

For the equilibrium:


"R_1+R_2=W+F_L"

"R_2L=W(0.5L)+F_L(d)"

"10R_2=5000(5)+15000(2)"

"R_2=5500\ N"

Thus,


"R_1+5500=5000+15000"

"R_1=14500\ N"